# On a novel UCP result and its application to inverse conductive scattering

### v2

@@ Abstract.
In this paper, we derive a novel Unique Continuation Principle (UCP) for a system of second-order elliptic PDEs system and apply it to investigate inverse problems in conductive scattering. The UCP relaxes the typical assumptions imposed on the domain or boundary with certain interior transmission conditions. This is motivated by the study of the associated inverse scattering problem and enables us to establish several novel unique identifiability results for the determination of generalized conductive scatterers using a single far-field pattern, significantly extending the results in [15, 21]. A key technical advancement in our work is the combination of Complex Geometric Optics (CGO) techniques from [15, 21] with the Fourier expansion method to microlocally analyze corner singularities and their implications for inverse problems. We believe that the methods developed can have broader applications in other contexts.
@@ 1. Introduction
where $-\pi\leq\vartheta_{m}<\vartheta_{M}<\pi$, $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$, and $\mathrm{i}=\sqrt{-1}$. Additionally, the two boundary lines of $\mathcal{S}$ are given by
@@
Theorem 1.1 shows a novel type of UCP result associated with a PDE system described by (1.3). The detailed proof of Theorem 1.1 is given in Section 2.
@@
Let $D$ and $\Omega$ be two bounded Lipschitz domains in $\mathbb{R}^{2}$ with connected complements $\mathbb{R}^{2}\setminus\overline{D}$ and $\mathbb{R}^{2}\setminus\overline{\Omega}$ respectively, where $\Omega\subset D$. Assume $\gamma_{1}\in\mathbb{R}_{+}$ and $\gamma_{2}\in L^{\infty}(D)\cap H^{2}(\Omega)$ with ${\rm supp}(\gamma_{2}-\gamma_{1})\subset\Omega$ are given. Suppose that $\Omega\cap B_{r_{0}}={\mathcal{S}}_{r_{0}}$, where ${\mathcal{S}}_{r_{0}}$ is defined by (1.2), $r_{0}\in\mathbb{R}_{+}$ is sufficiently small and $\mathbf{0}\in\partial\Omega$. The line segements $\Gamma^{\pm}_{r_{0}}$ are defined in (1.2). Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDE system:
@@ 1.2. The conductive medium scattering and invisibility
On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDE system (1.11). If there exists a nontrivial pair of solutions $(w,v)$ to (1.11) associated with $k$, then $k$ is referred to as a transmission eigenvalue and $(w,v)$ is the corresponding pair of transmission eigenfunctions of (1.11). Here, we would like to emphasise that throughout the article, $\eta$ can be taken to be identically zero, which reduces to the case with a normal scattering inhomogeneity of refractive index $q$, and all the results still hold equally. Hence, $(\Omega;q,\eta)$ is actually a generalized conductive medium scatterer and the transmission eigenvalue problem (1.11) is generalized transmission eigenvalue problem in our study, but we shall not distinguish it.
@@ 1.3. Connections to existing results and main contributions
For the main contribution a), in contrast to the existing weak, strong, or Cauchy-type UCPs for elliptic systems in the literature, we consider the UCP for the case where the components of the solution to (1.3) fulfill a transmission boundary condition on two intersecting line segments within the domain, forming an irrational corner. This transmission boundary condition arises from the invisibility or transparency phenomena in conductive medium scattering, which holds both practical and physical significance. We will discuss this aspect in more detail in Section 2. To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDE systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
@@
For the main contribution c), we can uniquely determine the refractive index in linear polynomial form and the constant conductive boundary parameter of the medium scatterer, whereas in [15] both the refractive index and the conductive parameter are constants. Moreover, we can relax the admissibility conditions in [15], which requires the non-vanishing of the total wave field at the corner points, to requiring the non-vanishing of either the total wave field or the gradient of the total wave field. In fact, the complexity inherent in the conductive polygonal-nest and polygonal-cell medium scatterer necessitates a detailed study of the singular behavior of the solution to a coupled conductive transmission PDE problem near a corner; see Lemma 3.2. Specifically, given that the refractive index is assumed to be a linear polynomial, it becomes crucial to demonstrate the Hölder regularity of the gradient of the total wave as derived from the direct problem (1.7), particularly near corners; see Lemma 3.1. With this regularity established, we propose another admissible condition, which describes the non-vanishing property of the gradient of the total wave field at the corner point. This admissible condition is generic, which can be fulfilled in more general physical scenarios; see more detailed discussions in Remarks 3.1 and 3.2. Under this generic admissible condition, we can then proceed recursively to determine the interior polygonal-nest structure.
@@
In this section, we aim to provide the proofs of Theorems 1.1-1.3. Specifically, we present a novel UCP result for an elliptic second-order PDE system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement. To prove Theorem 1.1, we first introduce some auxiliary lemmas. Drawing upon the results in [12, 41], we present the following lemma, which constructs a complex geometrical optics (CGO) solution.
@@
[18, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $\Delta+\gamma_{1}v=0$ in $\mathcal{S}_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$ and $\mathcal{S}_{r_{0}}$ is defined in (1.2). If $v$ is analytic in $\mathcal{S}_{r_{0}}$, then $v$ has the following spherical wave expansion in polar coordinates around the origin:
@@
$\lesssim\tau^{-\frac{13}{6}}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
@@
$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}.\end{matrix}\right)$
@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
In order to introduce the admissible class for studying the inverse problems, we need to study the Hölder-regularity of the total wave field $u$ which is associated with the scattering problem (1.7) near the corner $\mathcal{S}_{r_{0}}$. Then we have the following lemma, which shows that there holds that $u^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$.
@@
Let the corner $\mathcal{S}_{2r_{0}}=\mathcal{S}\cap B_{2r_{0}}$, where $\mathcal{S}$ is given by (1.1) and $B_{2r_{0}}$ is a disk centered at the origin with the radius $2r_{0}$ ($r_{0}\in\mathbb{R}_{+}$). Denote $\Gamma^{\pm}_{2r_{0}}=\Gamma^{\pm}\cap B_{2r_{0}}$, where $\Gamma^{\pm}$ is the boundary line of $\mathcal{S}$. Suppose that $u\in H^{1}(B_{2r_{0}})$ with $\ u^{+}=u|_{B_{2r_{0}}\setminus\overline{\mathcal{S}_{2r_{0}}}},\ u^{-}=u|_{\mathcal{S}_{2r_{0}}}$ satisfies that
@@
where $k$ is a positive constant and $q$ is in the form of (3.1). Assume that $u^{-}$ and $u^{+}$ are real analytic in $\mathcal{S}_{2r_{0}}$ and $B_{2r_{0}}\setminus\overline{\mathcal{S}_{2r_{0}}}$, respectively. Then there exist $\alpha\in(0,1)$ such that $u^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}).$
@@
Denote $\overline{u}$ by the analytic extension of $u^{+}|_{B_{2r_{0}}\setminus\overline{\mathcal{S}}_{2r_{0}}}$ in $B_{2r_{0}}$ since it is real analytic in $B_{2r_{0}}\setminus\overline{\mathcal{S}}_{2r_{0}}$. By virtue of the boundary condition of (3.7), one has $\overline{u}=u^{-}=u^{+}\ \mbox{on}\ \Gamma_{2r_{0}}^{\pm}.$
Let $w=u^{-}-\overline{u},$ then from (3.7) it can be directly calculated that
@@
        $\nabla u(\mathbf{x}_{p})\neq 0,\ u(\mathbf{x}_{p})=0,$ (3.9)
      where $\mathbf{x}_{p}\in{\mathcal{V}}(\Sigma_{i})$ and ${\mathcal{V}}(\Sigma_{i})=\{\mathbf{x}_{p,i}\}_{p=1}^{\ell_{i}}(3\leq\ell_{i}\in\mathbb{Z}_{+})$ is the vertex set of the convex polygon $\Sigma_{i}$. Here, $\{\Sigma_{i}\}_{i=1}^{N}$ is the polygonal-nest structure of $\Omega$ described by Definition 3.1.
@@
      If the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Sigma_{i}$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
@@
      Similarly, if the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Omega$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
@@
In previous works such as [21, 15], the admissible condition (3.8) plays a crucial role in deriving corresponding uniqueness results. However, in this paper, when (3.8) is violated, we propose another admissible condition, denoted as (3.9). There are three distinct cases for (3.9), characterized as follows:
@@
    $\partial_{1}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=\partial_{2}u(\mathbf{x}_{p})=0$,
@@
    $\partial_{2}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=\partial_{1}u(\mathbf{x}_{p})=0$,
@@
    $\partial_{1}u(\mathbf{x}_{p})\neq 0$, $\partial_{2}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=0$.
@@
$\begin{cases}\Delta u_{1}^{-}+k^{2}q_{1}u_{1}^{-}=0,\hskip 73.97733pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta u_{2}^{+}+k^{2}u_{2}^{+}=0,\hskip 83.93553pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ u_{1}^{-}=u_{2}^{+},\ \partial_{\nu}u_{1}^{-}=\partial_{\nu}u_{2}^{+}+\eta_{1}u_{2}^{+},\hskip 14.22636pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$
@@
$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}+\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.12)
According to Lemma 3.1, we know that $u_{2}^{+}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
$u_{2}^{+}=u_{2}^{+}(\mathbf{0})+\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2}+\delta u_{2}^{+},\ |\delta u_{2}^{+}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{+}\|_{C^{1,\alpha}}.$ (3.13)
@@
Case (a): $u_{2}^{+}(\mathbf{0})\not=0$. Combing (3.13) with (3.12), we can deduce that
$\eta_{1}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{4}I_{i},$ (3.14)
@@
$I_{1}$ $=-\eta_{1}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,\ I_{2}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}$
$I_{3}$ $=-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2}+\delta u_{2}^{+})u_{0}\mathrm{d}\sigma,$
@@
$|\eta_{1}u_{2}^{+}(\mathbf{0})|$ $\left|\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma\right|\gtrsim\frac{\Gamma(1)|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|}{|\tau e^{\mathrm{i}\varphi}|}-\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.17)
@@
$|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|=0.$

### v3

@@ Abstract.
In this paper, we derive a novel Unique Continuation Principle (UCP) for a system of second-order elliptic PDEs and apply it to investigate inverse problems in conductive scattering. The UCP relaxes the typical assumptions imposed on the domain or boundary with certain interior transmission conditions. This is motivated by the study of the associated inverse scattering problem and enables us to establish several novel unique identifiability results for the determination of generalized conductive scatterers using a single far-field pattern, significantly extending the results in [15, 21]. A key technical advancement in our work is the combination of Complex Geometric Optics (CGO) techniques from [15, 21] with the Fourier expansion method to microlocally analyze corner singularities and their implications for inverse problems. We believe that the methods developed can have broader applications in other contexts.
@@ 1. Introduction
where $-\pi\leq\vartheta_{m}<\vartheta_{M}<\pi$, $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\mathrm{i}=\sqrt{-1}$. Additionally, the two boundary lines of $\mathcal{S}$ are given by
@@
Theorem 1.1 shows a novel type of UCP result associated with a PDEs system described by (1.3). The detailed proof of Theorem 1.1 is given in Section 2.
@@
Let $D$ and $\Omega$ be two bounded Lipschitz domains in $\mathbb{R}^{2}$ with connected complements $\mathbb{R}^{2}\setminus\overline{D}$ and $\mathbb{R}^{2}\setminus\overline{\Omega}$ respectively, where $\Omega\subset D$. Assume $\gamma_{1}\in\mathbb{R}_{+}$ and $\gamma_{2}\in L^{\infty}(D)\cap H^{2}(\Omega)$ with ${\rm supp}(\gamma_{2}-\gamma_{1})\subset\Omega$ are given. Suppose that $\Omega\cap B_{r_{0}}={\mathcal{S}}_{r_{0}}$, where ${\mathcal{S}}_{r_{0}}$ is defined by (1.2), $r_{0}\in\mathbb{R}_{+}$ is sufficiently small and $\mathbf{0}\in\partial\Omega$. The line segements $\Gamma^{\pm}_{r_{0}}$ are defined in (1.2). Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDEs system:
@@ 1.2. The conductive medium scattering and invisibility
On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDEs system (1.11). If there exists a nontrivial pair of solutions $(w,v)$ to (1.11) associated with $k$, then $k$ is referred to as a transmission eigenvalue and $(w,v)$ is the corresponding pair of transmission eigenfunctions of (1.11). Here, we would like to emphasise that throughout the article, $\eta$ can be taken to be identically zero, which reduces to the case with a normal scattering inhomogeneity of refractive index $q$, and all the results still hold equally. Hence, $(\Omega;q,\eta)$ is actually a generalized conductive medium scatterer and the transmission eigenvalue problem (1.11) is generalized transmission eigenvalue problem in our study, but we shall not distinguish it.
@@ 1.3. Connections to existing results and main contributions
For the main contribution a), in contrast to the existing weak, strong, or Cauchy-type UCPs for elliptic systems in the literature, we consider the UCP for the case where the components of the solution to (1.3) fulfill a transmission boundary condition on two intersecting line segments within the domain, forming an irrational corner. This transmission boundary condition arises from the invisibility or transparency phenomena in conductive medium scattering, which holds both practical and physical significance. We will discuss this aspect in more detail in Section 2. To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDEs systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
@@
For the main contribution c), we establish a local uniqueness result for determining the shape of the polygonal-nest or polygonal-cell conductive medium scatterer. This is achieved by proposing a novel type of admissible condition: the non-anishing assumption of either the total wave field or its gradient. In contrast to the previous admissibility conditions in [15], which only required the non-vanishing of the total wave, the new conditions impose more requirements on the total wave. From the perspective of physical intuition, these new admissible conditions are more likely to be violated. As a result, the newly established uniqueness result is more applicable to general physical scenarios. For more in-depth discussions, please refer to Remarks 3.1 and 3.2. Furthermore, we can uniquely determine the refractive index in linear polynomial form and constant conductive boundary parameter of the medium scatterer, whereas in [15] both the refractive index and the conductive parameter are constants. In fact, the complexity inherent in the conductive polygonal-nest and polygonal-cell medium scatterer necessitates a detailed study of the singular behavior of the solution to a coupled conductive transmission PDE problem near a corner; see Lemma 3.2. We shall proceed recursively to identify the refractive index and constant conductive boundary parameter by leveraging the so-called complex geometrical optics (CGO) solution and the corresponding total wave field to give a dedicated microlocal analysis near the underlying corner.
@@
In this section, we aim to provide the proofs of Theorems 1.1-1.3. Specifically, we present a novel UCP result for an elliptic second-order PDEs system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement. To prove Theorem 1.1, we first introduce some auxiliary lemmas. Drawing upon the results in [12, 44], we present the following lemma, which constructs a complex geometrical optics (CGO) solution.
@@
[18, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $(\Delta+\gamma_{1})v=0$ in $\mathcal{S}_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$ and $\mathcal{S}_{r_{0}}$ is defined in (1.2). If $v$ is analytic in $\mathcal{S}_{r_{0}}$, then $v$ has the following spherical wave expansion in polar coordinates around the origin:
@@
$\lesssim\tau^{-\frac{13}{6}}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau})$
@@
$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}\end{matrix}\right).$
@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
In order to introduce the admissible class for studying the inverse problems of the unique identification of the shape, we need to study the Hölder-regularity of the solution of the coupled PDEs system (3.7) near the corner, where $u_{1}$ and $u_{2}$ are the total wave field corresponding to the two different scatterers $\Omega_{1}$ and $\Omega_{2}$, respectively. Indeed, the PDEs system (3.7) is derived from the proceed of proving the unique determination of the shape of the scatterer by contradiction (see (3.12) in the proof of Theorem 3.1).
@@
Let the corner $\mathcal{S}_{2r_{0}}=\mathcal{S}\cap B_{2r_{0}}$, where $\mathcal{S}$ is given by (1.1) and $B_{2r_{0}}$ is a disk centered at the origin with the radius $2r_{0}$ ($r_{0}\in\mathbb{R}_{+}$). Denote $\Gamma^{\pm}_{2r_{0}}=\Gamma^{\pm}\cap B_{2r_{0}}$, where $\Gamma^{\pm}$ is the boundary line of $\mathcal{S}$. Suppose that $u_{1},u_{2}\in H^{1}(\mathcal{S}_{2r_{0}})$ satisfy
@@
where $k$ is a positive constant and $q$ is in the form of (3.1). If $u_{2}$ is real analytic in $\mathcal{S}_{2r_{0}}$, then $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$.
@@
Let $w=u_{1}-u_{2},$ then from (3.7) it can be directly calculated that
@@
        if
        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0,\ \mbox{then}$
        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\nabla u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0,$ (3.9)
      where $m(B(\mathbf{x}_{p},\rho))$ is the measure of $B(\mathbf{x}_{p},\rho)$, $\mathbf{x}_{p}\in{\mathcal{V}}(\Sigma_{i})$ and ${\mathcal{V}}(\Sigma_{i})=\{\mathbf{x}_{p,i}\}_{p=1}^{\ell_{i}}(3\leq\ell_{i}\in\mathbb{Z}_{+})$ is the vertex set of the convex polygon $\Sigma_{i}$. Here, $\{\Sigma_{i}\}_{i=1}^{N}$ is the polygonal-nest structure of $\Omega$ described by Definition 3.1.
@@
      If the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(0,\pi)\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Sigma_{i}$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
@@
      Similarly, if the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(0,\pi)\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Omega$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
@@
In previous works such as [21, 15], the admissible condition (3.8) plays a crucial role in deriving corresponding uniqueness results. However, in this paper, when (3.8) is violated, we propose another admissible condition (3.9) to derive the unique determination of the shape of the conductive medium scatterer. There are three distinct cases for (3.9), characterized as follows:
@@
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
@@
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
@@
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$, $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$.
@@
$\begin{cases}\Delta u_{1}^{-}+k^{2}q_{1}u_{1}^{-}=0,\hskip 73.97733pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta u_{2}^{+}+k^{2}u_{2}^{+}=0,\hskip 83.93553pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ u_{1}^{-}=u_{2}^{+},\ \partial_{\nu}u_{1}^{-}=\partial_{\nu}u_{2}^{+}+\eta_{1}u_{2}^{+},\hskip 14.22636pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$ (3.12)
@@
$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}+\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.13)
According to Lemma 3.1, we know that $u_{1}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
$u_{1}^{-}=u_{1}^{-}(\mathbf{0})+\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2}+\delta u_{1}^{-},\ |\delta u_{1}^{-}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{+}\|_{C^{1,\alpha}}.$ (3.14)
@@
Case (a): $u_{1}^{-}(\mathbf{0})\not=0$. Combing (3.14) with (3.13), we can deduce that
$\eta_{1}u_{1}^{-}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{4}I_{i},$ (3.15)
@@
$I_{1}$ $=-\eta_{1}u_{1}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,\ I_{2}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}$
$I_{3}$ $=-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2}+\delta u_{1}^{-})u_{0}\mathrm{d}\sigma,$
@@
$|\eta_{1}u_{1}^{-}(\mathbf{0})|$ $\left|\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma\right|\gtrsim\frac{\Gamma(1)|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|}{|\tau e^{\mathrm{i}\varphi}|}-\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.18)
@@
$|\eta_{1}u_{1}^{-}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|=0.$